Find the electric flux through area of 2m^2 electric field of 200v/m. The angle between electric field and normal to area is 60°.


1
Expert's answer
2021-03-18T14:04:14-0400

By definition, the flux of the constant field throug some area is:


"\Phi =ES\cos \theta"

where "E = 200V\/m" is the electric field, "S = 2m^2" is the area, and "\theta = 60\degree" is the angle between electric field and normal to area. Thus, obtain:


"\Phi = 200\cdot 2\cdot \cos 60\degree = 200V\cdot m"

Answer. 200 V*m.